OpenStudy (anonymous):

how would I solve this? x2 - 4 = 0

OpenStudy (anonymous):

differnce of squares homie

OpenStudy (anonymous):

(x-2)(x+2)=0 x=2 and x=-2

OpenStudy (diyadiya):

x^2=4 x=+2 x=-2

OpenStudy (anonymous):

so every problem like this i have to put both the negative and positive?

OpenStudy (diyadiya):

if its square

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

it factors into (x+2)(x-2) by some rule. set it to 0 and solve. (x+2)(x-2)=0 x=2, -2

OpenStudy (anonymous):

some rule=zero product rule Get you rules straight

OpenStudy (anonymous):

x2 - 4 = 0 +4 + 4 x2 = 4 square both of them . to get x = the square root of 4 (: